OpenStudy (anonymous):

find the quotient 5 2/5 divided by 8 4/7 simplify your answer

OpenStudy (anonymous):

\[\frac{ 27 }{ 5 }\div \frac{ 60 }{ 7 }=\frac{ 27 }{ 5 }\times \frac{ 7 }{ 60 }\]

OpenStudy (anonymous):

\[\frac{ 9 }{ 5 }\times \frac{ 7 }{ 20 }\]

OpenStudy (anonymous):

So Its 5 * 20Then 100 Divided By 9 And 7

OpenStudy (anonymous):

\[\frac{ 9\times7 }{ 5\times20 }\]

OpenStudy (anonymous):

63 Over 100

OpenStudy (anonymous):

and that is simplified

OpenStudy (anonymous):

Thank You